Question Hep The data show the bug chirps per minute at different temperatures Find the regression equation letting the first variable be the independent [) variable Find the best predicted temperature for a time when a bug is chirping at the rate of 3000 chirps per minute Use a sigrificance level of 0.05 What is wrong with this predicted value? 1m7F) +005 1020부 1193 1243 995 1070 904 1193 1247 1075 p 81 82 73 872 883 843 Temperature ( F) What is the regression equation? (Round the xcoefficient to four decimal places an needed Round the constant to two decimal places as needed ) What is the best predicted temperature for a time when a bug ie chirping at the rate of 3000 chirps per minuse? The best predicted temperature when a bug is chirping at 3000 chirps per minute in (Round to one decimal place as needed) What is wrong with this predicted valus? Choose the correct answer below O A. it is only an approximation. An unrounded value would be considered accurate O B. It is unrealistically high. The value 3000 is far outside of the range of obaerved values Ο c. The first varable should have been the dependent vanable O D. Nothing ia wrong with this value. It can be treated as an accurate prediction. Cick to select your answer
Answer #1

data

chirp temperature
995 81
1070 82
904 73
1193 87.2
1247 88.3
1075 84.3

result Using Excel

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.95209559
R Square 0.906486012
Adjusted R Square 0.883107515
Standard Error 1.882682093
Observations 6
ANOVA
df SS MS F Significance F
Regression 1 137.4354 137.4354 38.77435 0.003387
Residual 4 14.17797 3.544492
Total 5 151.6133
Coefficients Standard Error t Stat P-value Lower 95%
Intercept 37.55386589 7.280152 5.15839 0.006705 17.34092
chirp 0.041714498 0.006699 6.226905 0.003387 0.023115

a)

y^ = 37.5539 + 0.0417 * x

=37.55 + 0.0417 *x

b)

when x = 3000

y^ = 37.5539 + 0.0417 * 3000

= 162.6539

=162.7

c)

option B) is correct